# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ 2.2.2 Alignment at multiple levels of abstraction
The Levels dataset allows to systematically study discrepancies between human and deep learning model decisions across different levels of complexity. Here, we demonstrate that soft-alignment can reduce those discrepancies: it makes it possible to incorporate the hierarchical structure of human conceptual knowledge into foundation model representations.
@@
Global coarse-grained. We found the largest differences between model and human similarity judgments before AligNet fine-tuning were present for global coarse-grained semantics. The base models achieved low odd-one-out accuracies of $36.09\%$ (ViT-B) -- $57.38\%$ (DINOv2 ViT-B). Soft-alignment significantly improved the alignment between human and model responses to the extent that all models performed above the human-to-human reliability score of $61.92\%$ (see Fig. 2d, leftmost column) with odd-one-out accuracies of $65.70\%$ (DINOv1 ViT-B) -- $68.56\%$ (DINOv2 ViT-B). The relative improvements in performance ranged from $19.48\%$ (DINOv2 ViT-B) -- $93.51\%$ (ViT-L).
@@
Local fine-grained. The performance of most base models (except for DINOv1 ViT-B) did not strongly correspond to human responses for fine-grained semantics either and, thus, model performances were far from the human noise ceiling of $65.92\%$. Prior to fine-tuning, models achieved poor alignment scores of $46.04\%$ (ViT-B) -- $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B which performed ($62.92\%$) significantly better than any other model in that setting (see Tab. 3 in the Supplementary Material). AligNet fine-tuning improved this mismatch to some degree but not as substantially as it did for the coarse-grained abstraction level. AligNet models achieved odd-one-out accuracies of $58.93$ (ViT-S) -- $62.92\%$ (DINOv1 ViT-B). This is equal to a relative increase in performance of $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L) (see Fig. 2d, center column).
Class-boundary. Supervised classifiers and image/text contrastive models performed close to the noise ceiling for class-boundary triplets prior to any fine-tuning. Their odd-one-out accuracies ranged from $81.96\%$ (SigLIP ViT-B) to $93.67\%$ (ViT-L) (see Tab. 3). However, the caption generator model (CapPa) often responded quite differently from humans in that setting and achieved a low odd-one-out accuracy score of $70.37\%$. AligNet fine-tuning changed the representation spaces of all models that we considered to be equally well aligned. There are small differences between the models but they are not statistically significant. Surprisingly, AligNet models’ odd-one-out accuracies were higher than the human noise ceiling of $89.21\%$ (see Fig. 2d, rightmost column) with the best model achieving an alignment score of $93.24\%$ (ViT-L). This means that the responses of AligNet models were more similar to the average human responses—since each triplet response is the majority response of the subject population—than the level of agreement among the human subjects themselves. The variance in accuracy among AligNet models was much smaller compared to the other two abstraction settings, with odd-one-out accuracies of $93.09\%-94.24\%$. The relative increases in performance for this setting were between $0.62\%$ (ViT-L) - $32.29\%$ (CapPa ViT-B). For more details about model performances on Levels see Tab. 3 in the Supplementary Material.
Human alignment depends on the abstraction level. We found that the best model before soft-alignment at a particular abstraction level was often not the best model after soft-alignment. While DINOv2 was the best aligned model before AligNet fine-tuning for the fine-grained (57.38%) and coarse-grained (57.72%) abstraction settings (see Tab. 3 in the Supplementary Material), it remained the best aligned model only for the fine-grained setting (62.24%) but did not improve as significantly as the other models for the coarse-grained (relative improvement of 18.78%) and class-boundary settings (relative improvement of 4.13%). On the other hand, ViT-L was the worst-aligned model before fine-tuning for the coarse-grained setting (35.45%), but after AligNet fine-tuning, it became the best-aligned model, achieving the highest odd-one-out accuracy (68.60%) across all models. The relative improvement in performance was 93.51%. In addition, ViT-L was both the best aligned model before (93.67%) and after (94.24%) AligNet fine-tuning for the class-boundary setting (both scores are higher than the human reliability score of 89.21%).
We observed a similar phenomenon for the coarse- and fine-grained human responses datasets of Peterson et al. [2016, 2018]. DINOv2 showed the highest Spearman rank correlation coefficient with the human similarity judgments prior to any fine-tuning for both abstraction levels but remained the best aligned model only for the fine-grained datasets after fine-tuning (see Tab. 2 in the Supplementary Material for details). This suggests that DINOv2 fine-tuned on AligNet best captured fine-grained human semantic structure whereas image/text contrastive and supervised models best reflected coarse-grained and class-boundary semantic structure respectively after the fine-tuning process.
@@
Model uncertainties correspond to human latencies after soft-alignment. In addition to discrete odd-one-out choices, we collected (continuous) human response times (RTs) for each participant and triplet. Human response latencies are indicative of the uncertainty of a participant’s choice (see Sec. 2.2.1). Thus, here we use them as a proxy of the participants’ uncertainty [Ratcliff, 1978; Kiani et al., 2014]. We measured the uncertainty of model responses as the discrete Shannon entropy of the three pairwise similarities within each triplet. The triplet uncertainties of the base models were not correlated with the human latencies for the coarse-grained ($\rho=-0.014-0.184$) and fine-grained settings ($\rho=0.047-0.160$) and showed a medium positive correlation for the class boundary setting ($\rho=0.208-0.432)$. After soft-alignment, the correlations substantially increased for all models and abstraction settings (see Fig. 2e). The increase in correlation was particularly striking for the coarse-grained abstraction level: AligNet models showed weak positive correlations for the fine-grained ($\rho=0.171-0.216)$, medium positive correlations for the coarse-grained ($\rho=0.331-0.376)$, and strong positive correlations for the class-boundary setting ($\rho=0.479-0.506)$. Correlations were measured as Spearman rank correlation coefficients.
@@ 2.3 Aligned models reflect the conceptual hierarchy
Figure 3: Aligned models reflect the semantic hierarchy. a: Before alignment, models trained with different losses have dissimilar representation structures—particularly those trained for supervised classification. After alignment, however, model representation structures are much more similar to each other. b: To understand the alignment, we study how the models’ representations change across the semantic hierarchy, from relations between images within the same subordinate category to relations across superordinate categories. c: The cluster-driven triplet sampling tends to produce triplets where two images have a closer relation than the third. d: The result is that the relations between image representations change following the semantic hierarchy—images from the same basic or superordinate category tend to move closer together, while those from different superordinate categories move farther apart. e: Visualizing the distance changes in more detail, with some superordinate categories boxed on the diagonal and labelled with icons. (Panels d-e are for the representations of ViT-B.)
How do the model representations change after soft-alignment? In Fig. 3 we show that while the model representations are dissimilar before alignment (especially those trained on classification tasks), they become much more similar to each other after soft-alignment (see Sec. 4.4 for details).
One reason for this convergence is that the models align better with the human semantic category hierarchy [cf. Rosch et al., 1976]. The organization during alignment is driven by two factors: first, as shown in Section 2.2, the labels generated by the model will tend to reflect human judgments and uncertainty, especially in the relationships at more abstract levels of the category hierarchy. However, the clustering process that we use to sample the triplets also reflects this hierarchical structure. In particular, it tends to produce triplets where the similar pair of images come from the same subordinate- or basic-level category, while the odd-one-out comes from either a different basic-level or superordinate category (Fig. 3c). Thus, our soft-alignment procedure embeds the global structure in multiple ways: both in clustering and labeling.
@@
The impact of this hierarchical alignment is that the relationships between representations of different images change depending on the semantic relationship between them. The representations of images from the same basic category tend to move much closer together, those of images from the same superordinate category tend to move somewhat closer together, and those from different superordinate categories tend to move farther apart. Thus, the alignment procedure drives model representations to reorganize according to the global structure of human knowledge, as we intended. (Results in Fig. 3d-e are for ViT-B; for other models see SI B.3.2.)
@@
As an illustrative example, in a standard ViT-B classifier the representations of lizards are initially closer to those of some plants and fruits due their similarity in texture, color, or background conditions; after alignment, they are naturally more similar to the representations of other animals and more distant from those of other, unrelated superordinate categories (see Fig. 1d). This organization will yield better generalization when lizards are anomalous among a category of fruits, or when a lizard image is used as one of the examples depicting a more abstract category like animals.
@@ 2.4 Alignment improves generalization and out-of-distribution robustness
What are the consequences of more human-like structured model representations induced by soft-alignment for machine learning tasks? To address this question, we investigated how alignment alters generalization and out-of-distribution robustness of foundation models across different machine learning downstream tasks.
To obtain a measure that systematically evaluates the quality of the models’ representations, we froze the weights of the neural network model and trained a linear classifier on top of the model’s image representations rather than training/fine-tuning the entire model. A key question for any real-world application is how well a model generalizes to new settings, tasks, and objects. Hence, we studied (a) one-shot classification tasks, (b) distribution shift and (c) out-of-distribution robustness.
One-shot classification. One-shot classification tasks are among the hardest scenarios for testing generalization: the objective is to classify images given only a single labelled example per class. In Fig. 4 we show one-shot performance of all models before and after soft-alignment on ten image-classification datasets from a wide range of domains, such as fine-grained bird types classification [Birds; Wah et al., 2011], land use classification from satellite imagery [UC Merced; Yang and Newsam, 2010], and colorectal cancer histology [Colon; Kather et al., 2016]. While for eight combinations of models and datasets, we find a small decrease in performance, the majority of cases (32) show an improvement, sometimes by a substantial margin (e.g., the accuracy of DINOv2 on the Pets dataset shows a 2.7 fold increase). Overall our alignment framework increased the generalization performance on these task significantly (p < 0.05). These results corroborate the hypothesis that a good coarse-level (re-)structuring of the representation space supports generalization in general and especially for tasks where little labelled data is available. In Appx. B.2.1 we show additional few-shot results for more tasks and models, and demonstrate that our method can be used in conjunction with other approaches that were specifically designed for improving the performance of pretrained foundation models in few-shot learning tasks. Often a caveat with those methods is that their application is restricted to image/text models because they require text representations [e.g., Zhang et al., 2022]. In contrast, soft-alignment can be applied to a model irrespective of how it was pretained.
@@
Distribution shift. A long-standing problem of applying machine learning algorithms in practice is distribution shift and domain adaptation [cf., Sugiyama et al., 2007; Sugiyama and Kawanabe, 2012; Farahani et al., 2021]: as machine learning systems are deployed, the data they were originally trained on often differs in subtle ways from the data they are being applied to. Models tend to fail in unexpected ways when the distribution of the data they should predict differs from that of the training data.
To measure whether the global category structure induced by fine-tuning on AligNet helps address these distributional shifts, we use the BREEDS benchmarks [Santurkar et al., 2020]. These were specifically designed explore generalization under input distribution shifts, by constructing datasets where the samples from training and test are sampled from different subpopulations. Fig. 4b shows the result of training a linear readout layer on the four BREEDS training sets, and evaluating them on the corresponding test sets. AligNet fine-tuning consistently improves performance across all benchmarks and models types. The image classifier (ViT-B) benefits the most with a more than twofold increase for the “entity13” benchmark.
@@
A further benefit of alignment is that it has a positive influence on the robustness of a model. Hendrycks et al. [2021b] collected a challenging dataset of natural images that are adversarial, in the sense that neural networks tend to consistently misclassify them, while they are obvious for humans. Again, we find that AligNet fine-tuning improves the accuracy across all VFMs that we investigate (cf. Fig. 4c); specifically, we observe improvements of up to $9.5$ percentage points, which corresponds to a $1.6$ fold improvement for the CapPa model.
Together, all of these results corroborate that aligning human and neural network model representations helps to generalize better, transfer to new tasks and data, and enhances model robustness. In other words, alignment is helpful to practically improve deep learning.
@@ 3 Discussion
The differences between natural intelligence and the capabilities of neural networks are the subject of long-standing debates [Fodor and Pylyshyn, 1988; Lake et al., 2017]. Despite the dramatic progress in AI of the past decade, these discussions persist, because even the most successful deep learning systems seem to fail in non-human-like ways [Lake et al., 2015; Lapuschkin et al., 2019; Geirhos et al., 2020; Thrush et al., 2022; Muttenthaler et al., 2023a; Lake and Baroni, 2023; Patraucean et al., 2023].
@@
In this work, we have highlighted—and addressed—a key deficiency in a broad class of vision foundation models: their representations do not adequately represent the multi-level conceptual structure of semantic knowledge that is natural to humans (see Sec. 2.2). We demonstrate this deficiency by collecting a new dataset of stratified human similarity judgments across multiple levels of abstraction—which we call Levels (see Sec. 2.2.2). To address this deficiency, we established a methodological framework for aligning deep learning models with human similarity judgments to build systems that exhibit more human-like behavior (see Sec. 2.1). By bootstrapping from a relatively small quantity of human data, we develop a surrogate neural network model of human object similarity judgments. Analogous to prior work on semi-supervised classification [e.g. Lee et al., 2013; Grandvalet and Bengio, 2004], we use this surrogate model to label a larger dataset for which we do not have human judgments. Specifically, we propose a cluster-based triplet-sampling technique for ImageNet images, which we then label with human-like odd-one-out responses from our surrogate model to create a large new synthetic dataset called AligNet.
@@
We fine-tuned a wide range of vision foundation models on AligNet, using a novel KL-divergence-based objective function for injecting human-like similarity structure (i.e., hierarchical conceptual knowledge) into the models. We systematically demonstrate that using this framework yields significantly increased alignment to human judgments on many cognitive science tasks (see Sec. 2.2), as well as better performance on various representative machine learning tasks (see Sec. 2.4). In other words, soft-alignment helps to alleviate the brittleness of machine learning models under changing environments. Moreover, our results illustrate how the broader paradigm of studying representational alignment [Kriegeskorte et al., 2008a; Sucholutsky et al., 2023] can not only yield insights by measuring how systems differ, but can be leveraged to actively align model representations with human conceptual hierarchies (see Sec. 2.3) to improve the models’ generalization and robustness.
@@
These results contribute to the long-standing debate over whether neural network models can capture the full range of human intelligence [Fodor and Pylyshyn, 1988; Marcus, 1998; McClelland et al., 2010; Holyoak and Hummel, 2014; Tenenbaum et al., 2011; Lake et al., 2017]. In particular, relational understanding is thought to be a distinguishing feature of human intelligence [Gentner, 2003; Waltz et al., 1999; Penn et al., 2008], and one relevant line of critique has argued that neural networks lack the capability to appropriately represent abstract relations like same and different [Marcus, 1998; Holyoak and Hummel, 2014], or to organize knowledge into hierarchies of concepts and their relations [Tenenbaum et al., 2011]. While some aspects of these critiques have been refuted in simpler synthetic settings [e.g. Geiger et al., 2023], similar criticisms persist for modern foundation models—including the specific claim that empirical weaknesses in vision models make them deeply flawed as cognitive models [Bowers et al., 2022]. Our results show that, while standard objectives do not adequately capture the hierarchy of relations within and across categories, these relations can be distilled into the models—and this procedure substantially improves the models’ resilience under the kinds of distribution shifts that have concerned some prior critiques.
@@
Although in this work we focused on the visual domain, similar global misalignments likely arise in other areas of machine learning research. For instance, in natural language processing, models are similarly trained with objectives that focus on distinguishing between close matches (e.g., prediction objectives that primarily distinguish words that occur in the current context from those that would occur in slightly different contexts, rather than considering their relations to less probable concepts). Applying alignment techniques to natural language processing may therefore analogously help to better capture the global structure of semantic and syntactic relationships among language inputs that these objectives miss.
Reconciling global and local structures is a long-standing problem across the natural sciences. For example, the inherent complexities of molecules and materials need both short-range (local) and long-range (global) components. The Schroedinger equation accurately describes this intricate interplay. Thus, modelling in quantum mechanics can use the laws of physics as inductive biases (see e.g. physics informed learning). Constraining [e.g., Chmiela et al., 2017, 2018; Schmitz et al., 2022] or regularizing [e.g., Karniadakis et al., 2021] models using physical priors/laws can explicitly include such global and local dependency components.
In contrast, the fields of vision and language processing do not offer equations that enable the direct access to local and global relationships between objects, scenes, or words. Therefore, in the AligNet framework we propose to use human knowledge as a regularizer to achieve the desired representations. A more general insight is that the implicit knowledge of humans can be made readily accessible to learning systems via soft-alignment. Thus, learning systems become more human-like because in addition to local information they can now reflect global relationships. We hypothesize that this approach could also be applied to other domains where human judgements of global relations are accessible, but equations are not.
@@
We remark that AI systems have been successfully adopted in many areas across the sciences and industries. However, these deployments lead to many practical [Paleyes et al., 2022] and conceptual [Amodei et al., 2016] concerns such as trustworthiness, biases, safety issues, and miscalibration. It is therefore increasingly important to identify the reasons why these systems occasionally fail (e.g., global misalignment, non-robustness, spurious correlations, or shortcut learning) and how to alleviate these failures. Our work advances the understanding of the deficiencies of a major class of current systems, and simultaneously shows a viable path for ameliorating these deficiencies by alignment with human judgments.
@@
In order to most effectively build on this work, there are a number of limitations that could be addressed in future efforts. First, the neural network models we used neither account for context in similarity judgments nor higher-order relations. These factors could possibly be accounted for with more complex models, such as contemporary vision-language models [e.g. Team et al., 2023; OpenAI et al., 2023], with which performance and alignment across different contexts could be readily explored. In this work we generally treat the human representation as if it were common among all humans, modulo some intrinsic uncertainty across individuals. However, human representations may vary systematically across individuals, cultures, and so on. Exploring this dimension of variation in human judgments would be an important topic for future investigation. Finally, human judgment is full of flaws, intrinsic contradictions and discrepancies. Given these issues, perfect alignment to human performance may not always be desirable for a technical system. Thus, future work could explore how to best learn from human knowledge without adopting human imperfections.
@@
In summary, we have provided an initial blueprint of how to distill global, human-like similarity structures into the representations of modern deep neural networks. We demonstrated a viable path toward a best-of-both-worlds representation that is both more consistent with human cognition and more practically useful in the machine learning domain, paving the way toward more robust, interpretable, and human-like artificial intelligence systems. We hope that our work will inspire more general approaches to (softly) aligning models by distilling priors from human cognition into their representations.

### v4

@@ 2.1.1 Alignment at multiple levels of abstraction
Levels dataset. Because prior cognitive datasets were not specifically targeted for assessing knowledge of vision foundation models across levels of abstraction, we collected a novel dataset of human judgments—which we call Levels—that is based on the triplet odd-one-out task, but stratified across different levels of the semantic hierarchy. Specifically, we collect global coarse-grained semantic, which require deciding on the odd-one-out among broadly different categories; local fine-grained semantic, involving discerning subtle distinctions within the same category; and class-boundary, testing the capacity to identify category boundaries. For details see Sec. 4.3.
Alignment at multiple levels. The levels dataset allows us to systematically study discrepancies between human and model decisions across these different levels. We find that our soft-alignment method reduces these discrepancies at all levels, but especially for the global coarse-grained judgments, as we hypothesized. Specifically:
@@
Global coarse-grained. This level shows the largest improvements. The base models achieved low accuracies of $36.09\%$ (ViT-B) - $57.38\%$ (DINOv2 ViT-B). AligNet models significantly improved; all models performed well, with accuracies of $65.70\%$ (DINOv1 ViT-B) - $68.56\%$ (DINOv2 ViT-B)— above the human-to-human reliability score of $61.92\%$ (see Fig.˜2d, leftmost column). That is, the AligNet models’ responses were more similar to average human responses (since each triplet response is the majority response of the subjects) than the level of agreement among the human participants. AligNet models’ relative improvements ranged from $19.48\%$ (DINOv2 ViT-B) - $93.51\%$ (ViT-L).
@@
Local fine-grained. Most base models did not strongly align to human responses for fine-grained semantics either; all models achieved poor alignment scores of $46.04\%$ (ViT-B) - $57.72\%$ (DINOv2 ViT-B), except for DINOv1 ViT-B which performed significantly better($62.92\%$; near the human noise ceiling of $65.92\%$; see SI Tab. 3). AligNet models achieved increased accuracies of $58.93$ (ViT-S) - $62.92\%$ (DINOv1 ViT-B); relative improvements $7.84\%$ (DINOv2 ViT-B) - $46.03\%$ (ViT-L) (see Fig.˜2d, center column).
Class-boundary. Supervised classifiers and image/text contrastive base models performed close to the noise ceiling; accuracies ranged from $81.96\%$ (SigLIP ViT-B) to $93.67\%$ (ViT-L). Others performed worse; the CapPa captioning model achieved $70.37\%$. AligNet fine-tuning brought all models to a similar level, achieving accuracies up to $93.24\%$ (ViT-L), higher than the human noise ceiling of $89.21\%$ (see Fig.˜2d, rightmost column). Relative improvements were $0.62\%$ (ViT-L) - $32.29\%$ (CapPa ViT-B). For more performance details see SI Tab. 3.
@@
AligNet model uncertainties correspond to human latencies. We also collected (continuous) human response times (RTs), which we use as a proxy of the participants’ uncertainty [67, 36]. We measured model uncertainty as the entropy of the three pairwise similarities within each triplet. Base model uncertainties were not correlated with human RTs for the coarse ($\rho=-0.014-0.184$) and fine-grained settings ($\rho=0.047-0.160$) and moderately correlated for the class boundary setting ($\rho=0.208-0.432)$. All AligNet models showed substantially increased uncertainty alignment across all levels (see Fig.˜2e), especially the coarse-grained abstraction level ($\rho=0.479-0.506)$.
Evaluating other model classes. To confirm that the aforementioned representational weaknesses are present in other classes of deep learning models, we evaluated two state-of-the-art natively-multimodal LLM, Gemini 2.0 Flash and Gemini 2.5 pro [24], on Levels. These VLMs perform similarly to—or slightly better than—the better pretrained vision models across all levels; however, they still substantially underperform our AligNet fine-tuned models (see SI B.1.7). Additionally, we evaluated models trained on Ecoset [51], an ecologically motivated natural image dataset, and find that Ecoset models have severe difficulties matching the human similarity judgments (see SI B.1.6). These results confirm that the AligNet fine-tuning offers greater improvements in human alignment than merely incorporating language modeling or ecological training.
@@ 2.2 Aligned models reflect the conceptual hierarchy
Figure 3: Aligned models reflect the semantic hierarchy. a: Before alignment, models trained with different losses have dissimilar representation structures—particularly those trained for supervised classification. After alignment, however, model representation structures are much more similar to each other. b: To understand the alignment, we study how the models’ representations change across the semantic hierarchy, from relations between images within the same subordinate category to relations across superordinate categories. c: The cluster-driven triplet sampling tends to produce triplets where two images have a closer relation than the third. d: The result is that the relations between image representations change following the semantic hierarchy—images from the same subordinate, basic, or superordinate category tend to move closer together, while those from different superordinate categories move farther apart. e: Visualizing the distance changes in more detail, with some superordinate categories boxed on the diagonal and labelled with icons. (Panels d-e are for the representations of ViT-B.)
@@
How do model representations change after soft-alignment? In Fig.˜3 we show that while the model representations are dissimilar before alignment, they become more aligned with each other after soft-alignment. This convergence is driven by models aligning better with the human conceptual hierarchy [Fig.˜3b; cf. 72]. Our soft-alignment procedure embeds this global structure at two levels: first in the cluster-based sampling (Fig.˜3c), and then in the soft-labels (see above). Because of these factors, the relationships between image representations change during alignment according to their semantic relationship; representations of images from the same basic category tend to move closer together, those of images from the same superordinate category tend to move somewhat closer, and those from different superordinate categories tend to move apart (all effects are highly-significant; $t$s>3.93>3.93, $p$s<0.001<0.001).
@@
As an illustrative example, in a base ViT-B the representations of lizards are close to those of some plants and fruits due their similarity in texture, color, or background; after alignment, they become similar to representations of other animals and more distant from those of other, unrelated superordinate categories. This reorganization yields better generalization, e.g. when a lizard image is used as an example depicting an abstract category like animals.
SI B.3.2 presents detailed analyses, including reorganization at high levels such as living vs. nonliving, across layers, and in other models. Furthermore, SI B.3.3 shows that where the AligNet model and a baseline unaligned model disagree, human judgments are strongly correlated with those of AligNet, but not the ablation model—in fact, emphevery human participant in the study agreed more with AligNet.
@@ 2.3 Alignment improves generalization and out-of-distribution robustness
How do human-aligned representations affect performance on machine learning tasks? We investigated how alignment improves generalization and out-of-distribution robustness across a variety of downstream tasks.
One-shot classification. We first test an extremely-challenging generalization setting: classifying images given only a single labelled example per class. In Fig.˜4 we show one-shot performance before and after soft-alignment on ten image-classification datasets from varied domains, such as fine-grained bird [Birds; 82] and flower [Flowers; 58] types classification, multi-domain natural image classification [ImageNet; 16], and scene recognition [Places365; 89]. The majority of cases (32 of 40) show an improvement, sometimes by a substantial margin (e.g. DINOv2 shows a $2.7\times$ increase on the Pets dataset); overall, alignment significantly increases the generalization performance on these tasks (p < 0.05). These results show how human-aligned representations support strong generalization from little data. See SI B.2.1 for additional results, including the complementary benefits of combining our method with other approaches to few-shot generalization.
Distribution shift. A long-standing problem for applying machine learning algorithms is distribution shift [cf., 76, 18]: in deployment, data often differ in subtle ways from the training data, leading to unexpected model failures. To evaluate whether the global structure induced by alignment helps ameliorate this issue, we evaluated our models on the BREEDS benchmarks [74]—which specifically tests generalization under input distribution shifts, using datasets where training and test data points are sampled from different subpopulations. Fig.˜4b shows that AligNet fine-tuning consistently improves performance significantly across all benchmarks and models types (especially the image classifier, ViT-B). See SI B.2.3 for further results.
@@
Alignment also improves robustness. We evaluate on ImageNet-A [32], a challenging dataset of natural images that are adversarial (i.e., models tend to misclassify them, but humans perform better). Again, alignment improves accuracy for all models (cf. Fig.˜4c), with improvements of up to $9.5$ percentage points ($1.6\times$ improvement for CapPa). While our method is not designed for improving OOD robustness, its improvements are comparable to state-of-the-art methods designed for precisely this problem (see Tab. 6 in the SI).
Togerther, these machine learning results corroborate that AligNet fine-tuning improves the generalization, transfer, and robustness of model representations.
@@ 3 Discussion
The differences between natural intelligence and the capabilities of neural networks are the subject of long-standing debates [19, 45]. Despite the dramatic recent progress in AI,these discussions persist, because deep learning systems still seem to fail in non-human-like ways [46, 23].
@@
Here, we have highlighted—and addressed—a key deficiency in a broad class of vision foundation models: their representations do not adequately represent the multi-level conceptual structure of human semantic knowledge (see Sec. 2.1). We demonstrate this deficiency through Levels, a new dataset of human similarity judgments across multiple levels of abstraction. To address this deficiency, we established a methodological framework for aligning deep learning models’ representations with human judgments to create more human-aligned systems. This framework involves bootstrapping from a small quantity of human data to train a surrogate teacher model, and using this teacher to create a large new synthetic dataset (AligNet), which we use to fine-tune various vision-foundation models to inject human-aligned structure
@@
This approach yields significantly increased alignment to human judgments on cognitive science tasks (Sec. 2.1), and better generalization and robustness on representative machine learning tasks (Sec. 2.3). Thus, soft-alignment helps to reduce the brittleness of machine learning models under changing environments. Moreover, our results illustrate how the broader paradigm of studying representational alignment [42, 75] can not only yield insights about how systems relate, but can be leveraged to actively align model representations with human knowledge to improve the models’ generalization.
@@
These results contribute to long-standing debates over which features of human intelligence neural networks may lack [19, 78, 45]. In particular, one line of critique argues that neural networks lack the capability to appropriately represent abstract relations like same and different [50, 33], or to organize knowledge into hierarchies of concepts [78]. While aspects of these critiques have been refuted in simple synthetic settings [e.g. 21], similar criticisms persist for modern foundation models[9]. Our results show that, while standard training objectives do not adequately capture hierarchical category relations, these relations can be distilled into the models—which improves the models’ resilience under the distribution shifts highlighted in prior critiques. These results show that hierarchical representations may emerge from a system that is neither explicitly hierarchically structured, nor trained explicitly on the hierarchy.
@@
Although we focused on vision, similar global misalignments likely arise in other areas of research. For instance, in natural language processing, models are similarly trained with objectives that focus on distinguishing between close matches (e.g., prediction objectives that primarily distinguish words that are likely to occur, rather than considering their relations to less probable concepts). Applying alignment techniques may therefore analogously help to better capture the global structure of semantic and syntactic relationships among language inputs that these objectives might miss.
@@
More broadly, AI systems have been successfully adopted in many areas. However, these deployments lead to practical [60] and conceptual [4] concerns about trustworthiness and safety. It is therefore increasingly important to identify the reasons why these systems occasionally fail and how to alleviate these failures. Our work advances the understanding of the deficiencies of vision model representations, and simultaneously shows a viable path for ameliorating these deficiencies by alignment with human judgments.
@@
Our work has several limitations that could be addressed in future efforts. First, the models we used neither account for context in similarity judgments nor higher-order relations. Second, human representations may vary systematically across individuals, cultures, and so on. Finally, human judgment is full of flaws, intrinsic contradictions and discrepancies. Given these issues, perfect alignment to human performance may not always be desirable for a technical system. Thus, future work could explore how to best learn from human knowledge without adopting human imperfections.
@@
In summary, we have provided an initial approach to distill global, human-aligned similarity structures into the representations of modern deep neural networks. We demonstrated an efficient path toward a best-of-both-worlds representation that is both more consistent with human judgments and more practically useful, paving the way toward more robust, interpretable, and human-aligned artificial intelligence systems. We hope that our work will inspire more general approaches to (softly) aligning models by distilling human priors into their representations.